Suppose that people in an economy want to hold for transaction purposes an amount of money equal to one-fourth of national income. The following table shows the combined amounts of money people want to hold for precautionary and speculative purposes.

INTEREST RATE
AMOUNT OF MONEY DEMANDED
12% $80
10 100
8 120
6 140
4 160
2 180


(a) If national income is $400 and the interest rate is 10 percent, what is the total amount of money people will want to hold?
(b) If national income is $800 and the supply of money is $340, what is the equilibrium interest rate?
Answer #1

We are given that people want to hold for transaction purposes an amount of money equal to one-fourth of national income

(a) If national income is $400, money demanded for transaction purpose is $100. At the interest rate of 10 percent, money demanded for precautionary and speculative purposes is 100. Hence, the total amount of money people will want to hold is 100 + 100 = $200.

(b) If national income is $800, then people want to hold for transaction purposes, 800 / 4 = $200. Money supply should be equal to money supply implying that

200 + money demanded for precautionary and speculative purposes = 340

money demanded for precautionary and speculative purposes = 140.

Hence, total money demanded is 200 + 140 = 340 and the supply of money is $340, which indicates from the table that the equilibrium interest rate is 6%.